Asset Publisher

Herschel Status Report - July/August 2010

Herschel Status Report - July/August 2010

Report for period 23 June to 23 August 2010Mission operations of the Herschel space observatory continued nominally during the reporting period, with the spacecraft and subsystems all performing as expected. The first Open Time (OT1) Call for non Key Programme observing proposals closed on 22 July 2010. A total of 576 proposals were received.

In mid-August, Herschel was exposed to the first solar proton storm since launch. The teams of SPIRE and PACS (which were the two active instruments around the time in question) were alerted of possible impacts, and the Data Processing Team at the Herschel Science Centre (HSC) is screening for potential effects on data quality (increased noise levels) in the observations performed during this period.

Spacecraft

Star tracker warm pixels
The "bad pixel" table has been successfully uploaded to the spacecraft on 14 July 2010. Herschel's attitude control and measurement system (ACMS) will now correct for the effect of identified "warm" pixels that might otherwise impact the accuracy of pointings and the spatial reconstruction of mapping observations.

Since the star tracker CCD temperature reduction to -10°C at the end of March 2010, there has been no recurrence of the "speedbumps" in scan maps which were the result of the "warm" pixels (see the May 2010 status report for a summary of this issue and the counter measures).

Based on recent data it is estimated that 3 new "warm" pixels are acquired per week. Extrapolating this figure suggests that the maximum number of entries the "bad pixel" table can hold (128) will be hit in the last quarter of 2010. When this limit is reached, the rules that define what constitutes a "warm" pixel will have to be adapted from absolute to relative criteria that allow ignoring the worst, instead of all, such blemishes. Because the CCD has a quarter of a million pixels there is no real reason for concern, but the situation continues to be monitored through CCD dumps which are performed every other week.

Payload

Except for the occasional single event upsets (SEUs), all three instruments continue to operate nominally in the Routine Science Phase.

In three recent PACS spectroscopy observations, some of the detector voltage supplies were reporting non-nominal values. A thorough investigation of historical data uncovered more occurrences of such apparent voltage changes. However, only four of these (the three recent ones, and one from a few months ago which had escaped detection due to the effects being very subtle) were found to have actually corresponded to real voltage changes on the detectors, with an impact on the quality of the science data returned. The four affected observations have been rescheduled.

In general this anomaly only affects reference voltage settings. In the few cases where actual detector voltages are impacted this is not an instrument safety issue, because of the very high impedance of the detectors.

The HIFI instrument continues to experience an occasional SEU in the Local Control Oscillator Unit. In the majority of these cases, science data are not affected, either because the instrument is not prime, or because non-critical memory is affected.

In the previous status report (June 2010), based on data available up to that point, the susceptibility of the 3 instruments to SEUs seemed to be widely different. These differences have now significantly decreased or even vanished. Work is under way to further minimise the operational impacts of the SEUs by defining or further refining detection and recovery procedures.

Ground Segment

Mission Operations
During the reporting period, mission operations have been conducted with the support of the ESA New Norcia ground station. Observational data stored on-board Herschel was received on ground during communication passes lasting approximately three hours.

Archiving
The ground segment is operating nominally. Data products are generated routinely and ingested into the Herschel Science Archive (HSA).

Future Milestones

  • 1 November 2010: Announcement of approved OT1 proposals


---
Legal disclaimer
This report is based on the Herschel mission manager's report dated 23 August 2010. Please see the copyright section of the legal disclaimer (linked from the home page http://sci.esa.int) for terms of use.

Last Update: 1 September 2019
23-Feb-2020 21:30 UT

ShortUrl Portlet

Shortcut URL

https://sci.esa.int/s/wQernXw

Images And Videos

Related Publications

Documentation